please answer all this question

thank you!BACK NEXT CALCULATOR MESSAGE MY INSTRUCTOR FULL SCREEN PRINTER VERSION For July: 1. The June 30 bank reconciliation indicated
Answer #1
a
Deposits in Transit at July 31 1431 =16660-(15970-741)
b
Outstanding checks at July 31 1781 =17710-(16640-711)
c
Deposits in Transit at August 31 3910 =2540+(26930-25560)
d
Outstanding checks at August 31 3410 =2170+(25200-23960)